OpenStudy (anonymous):

Determine whether the function is even, odd, or neither. f(x)= |t-3|

OpenStudy (anonymous):

Sorry it's G(t)= | t-3 |

OpenStudy (lgbasallote):

absolute values are usually symmetrical to the y-axis...does that help?

OpenStudy (anonymous):

Kind of

OpenStudy (lgbasallote):

wonderful

OpenStudy (lgbasallote):

btw...when i say symmetrical to the y axis...i mean the line of symmetry is along the y-axis

OpenStudy (anonymous):

usually... but that function G(t) = | t-3 | is not....

OpenStudy (anonymous):

Ok so the answer to the question is odd? Because its diiferent from the original equation?

OpenStudy (anonymous):

Because in class it said that you plug in (-t) to the equation and then you simplify and it got a different equation do is it odd?

OpenStudy (lgbasallote):

lol wow. who said that? G(-t) = G(t) <----even -G(t) = G(t) <----odd -G(-t) = G(t) <----symmetric to origin

OpenStudy (anonymous):

I'm confused then do how do you do it then?